OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

Parth (parthkohli):

Use the distance formula.

OpenStudy (anonymous):

what is it?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow d = \sqrt{(x_{2} - x_{1})^2 + (y_{2} - y_{1})^2} }\)

Parth (parthkohli):

here it will be: \(\Large \color{Black}{\Rightarrow d = \sqrt{(-1 - 8)^2 + (-1 - 11)^2} }\)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{81 + 144} }\) \(\Large \color{Black}{\Rightarrow \sqrt{225} }\) \(\Large \color{Black}{\Rightarrow 15 }\)

Parth (parthkohli):

15 units.

OpenStudy (anonymous):

thanks for breaking it down

Parth (parthkohli):

My pleasure girlly! :P i like your singing btw again

OpenStudy (anonymous):

thanks!